OpenStudy (anonymous):

why does 2 log9 turns into 3 log4

OpenStudy (anonymous):

sorry i meant turn into 4log 3

OpenStudy (jameshorton):

wat grade are u in

OpenStudy (irishboy123):

\(\large a \ log x = log \ x^a\) \(\large (a \times \ b) \ log \ x = \log \ x^{a \times b} = b \ log \ x^a = a \ log \ x^b\) so \(\large 2 \ log \ 9 = 2 \ log \ 3^2 = 4 \ log \ 3 \ \ [= log \ 3^4 = log \ 81]\)